OpenStudy (samanthacrump):

BIO EXPERTS! HELP!HARDY WEINBERG...... In a population of a plant species, 225 plants are tall, while 26 are short. The height of the plant is controlled by a single gene, and only two alleles are present in this population:T (tall) and t (short), with t being recessive to T. The table below shows the number of short and tall plants over the next several generations:

OpenStudy (samanthacrump):

Generation | Tall Plants (T) | Short Plants (t) |_____1____|______225_____|____26_____| |_____2____|______275_____|____28_____| |_____3____|______233_____|____23_____| |_____4____|______281_____|____21_____| |_____5____|______295_____|____18_____|

OpenStudy (samanthacrump):

1. Using the Hardy-Weinberg equation, is evolution occurring in this population? Explain your answer. 2. Describe how an increase in competition for sunlight would affect the frequency of T and t over time.

OpenStudy (tkhunny):

Are the allele and genotype ratios remaining relatively constant?

OpenStudy (samanthacrump):

relatively yes. there is no increasing or decreasing. right?

OpenStudy (samanthacrump):

i think.. maybe im wrong

OpenStudy (tkhunny):

Look at the ratios. Generation | Tall Plants (T) | Short Plants (t) |_____1____|______225_____|____26_____| 225/26 = 8.65 |_____2____|______275_____|____28_____| 275/28 = 9.82 |_____3____|______233_____|____23_____| 233/23 = 10.13 |_____4____|______281_____|____21_____| 281/21 = 13.38 |_____5____|______295_____|____18_____| 295/18 = 16.39 H/W tells us that there should be no significant trend, UNLESS there are evolutionary forces at play.

OpenStudy (samanthacrump):

so is there evolution occuring??

OpenStudy (tkhunny):

That is the assumption of the theory, based on the ratio trend. Of course, finding the actual influence might be a bit trickier. To DOUBLE in only five years is shocking. We should be able to find that source.

OpenStudy (tkhunny):

Important: If we find some cause, particularly man-made, fixing it should restore equilibrium in very few generations - maybe only one!